OpenStudy (anonymous):

In triangle ABC, BE = 4, EC = 10, AD = x, and DC = 11. Find x so that segment DE and AB are parallel.

OpenStudy (anonymous):

what is E?

OpenStudy (anonymous):

thats all i have for the equation

OpenStudy (anonymous):

i assume that D is a point on ac and e is a point on cb if so for de to be parallel to ab CE /EB = CD / DA SO PLUG THE 3 VALUES INTO THIS and u can calculate AD

OpenStudy (anonymous):

10/4 = 11/AD

OpenStudy (anonymous):

my choices are 3.6 4.4 40/11 3.5

OpenStudy (anonymous):

4.4

OpenStudy (anonymous):

AD = 11*4 / 10 = 4.4

OpenStudy (anonymous):

44/10=4.4